OpenStudy (anonymous):

What mass of copper is required to replace silver from 4g of silver nitrate dissolved in water? Cu(s) + AgNO3 -> Cu(NO30)2 + Ag

OpenStudy (rushwr):

We can do it up to a certain limit with the given information. As we can see in the equation the mole ratio of AgNO3: Ag is 1:1 so 1st we have to find the number of moles of Ag: we know that no. of moles(n) = mass(m) divided by molar mass(M). \[n=\frac{ m }{ M }\] So we can find no of moles = 4/107gmol^-1 hence we found that no. of mole of Ag= no. of moles of AgNO3 Using the same euqation we can find the mass m= n*M we know the molar mass of AgNO3 would be around 170gmol^-1 So we can find the mass of AhNO3 required

OpenStudy (rushwr):

I got the mass of AgNO3 as 6.355g But if they have given the molar masses different . Please substitute them and get the answer :)